The range of a projectile fired at angle 15degrees is 50m. If it is fired with same speed with an angle of 45degrees.

Grade:6

1 Answers

Ghufran Amanullah
7 years ago
First we find the velocity by given dataAngle=15° ,Range=50 m, taking g=9.81 m/sR=Vo^2*sin2(angle)/g50=Vo^2*sin2(15)/9.8150=Vo^2*0.5/9.81 (since sin(30)= 0.5)490.5=Vo^2*0.5 (9.81*50=490.5)980=Vo^2Hence Vo=31.31Now again subsituting Vo=31.31, angle=45° and g=9.81R=(31.31)^2*sin2(45)/9.81R=980*1/9.81 (since sin(90)=1)R=100 mThe final result is R=100 m